# MIRROR: Multi-Modal Pathological Self-Supervised Representation Learning via Modality Alignment and Retention

### v2

@@ III-E Style Clustering Module
III-F Global Optimization Objective
The global optimization objective is formally defined as:
@@ IV-A Datasets
Four distinct cohorts from the publicly available TCGA dataset [17] were utilized for evaluation: Breast Invasive Carcinoma (BRCA), Non-Small Cell Lung Cancer (NSCLC), Renal Cell Carcinoma (RCC), and Colon and Rectal Adenocarcinoma (COADREAD). Each cohort comprises specific subtypes, as detailed in Table III.
@@
WSIs obtained from these cohorts were processed into patches at 20×20$\times$20 × magnification, and feature representations were extracted using ResNet-50 [53] and Phikon [54]. Corresponding transcriptomics data were collected from Xena [55] and preprocessed using the proposed novel pipeline, reducing the original 198,620 genes to 40,146 for the BRCA cohort, 10,234 for the NSCLC cohort, 10,303 for the RCC cohort, and 20,056 for the COADREAD cohort, forming the novel refined transcriptomics datasets.
@@ IV-B1 Data Preprocessing
For WSIs, adhering to the conventional preprocessing protocol established in [56], the foreground tissue regions are first segmented using the Otsu’s method [57]. These segmented regions are then divided into patches, forming an instance bag that is processed by a pre-trained patch encoder to extract feature representations. To enable batched training, we perform random sampling to select a fixed number of feature representations, employing replacement if the available number of representations is less than the required count.
@@ IV-B2 Experiment Setup
Our experiments evaluate the efficacy of the proposed model, MIRROR, on cancer subtyping and survival analysis tasks by concatenating the outputs from the two modality-specific encoders and feeding the combined vector into a linear classifier. The evaluations utilize few-shot learning and linear probing with 5-fold cross-validation. To assess generalization ability, two different pre-trained patch encoders are used for evaluation.
The model is optimized using the Adam optimizer [58] with a learning rate of $2\times 10^{-5}$. Implementation is conducted in PyTorch [59], and the model is trained for 100 epochs with a batch size of $16$. All training sessions are performed on a single NVIDIA GeForce RTX 3090 GPU paired with an Intel i7-12700K CPU and 32GB of memory.
@@ IV-C Downstream Tasks
Two tasks are used for downstream evaluations, namely subtype classification and survival analysis. We compared the proposed MIRROR against ABMIL [60], PORPOISE [61], Linear Classifier, and TANGLE [12], utilizing two distinct backbones under both 10-shot and linear probing settings. Among these methods, ABMIL and Linear Classifier take histopathology as input, whereas the other models integrate both histopathology and transcriptomics data. To ensure fair comparisons and optimal performance, all methods are trained on the proposed transcriptomic datasets. Notably, the original TANGLE framework evaluates only histopathology features. To comprehensively assess its performance and maintain methodological consistency, we extend TANGLE by incorporating both histopathology and transcriptomic features, aligning its input with that of MIRROR.
@@ IV-C1 Subtype Classification
As illustrated in Table I, the proposed MIRROR consistently achieves superior performance, demonstrating its effectiveness in integrating histopathology and transcriptomic data. On the TCGA-NSCLC dataset, MIRROR achieves the highest accuracy and F1 score across all settings, outperforming all baseline methods. Notably, its strong performance remains consistent regardless of the chosen backbone architecture, highlighting its robustness on the TCGA-NSCLC dataset. Similarly, on the TCGA-BRCA dataset, MIRROR achieves state-of-the-art results, surpassing the second best model’s (TANGLE) result by 1.5% in F1 score under linear probing setting with the Phikon backbone. Additionally, MIRROR significantly outperforms TANGLE by 6.4% in F1 score on the TCGA-RCC dataset under the 10-shot setting with the Phikon backbone. On the TCGA-COADREAD dataset, it surpasses TANGLE by 1.3% in accuracy under the linear probing setting with the Phikon backbone. These findings validate the robustness and generalizability of MIRROR across different cancer subtypes and backbone architectures.
@@ IV-C2 Survival Prediction
As shown in Table II, the proposed MIRROR achieves significant performance improvements over all baseline methods. On the TCGA-NSCLC dataset, MIRROR surpasses PORPOISE by $4.0\%$ and achieves a $7.6\%$ improvement compared to ABMIL under the 10-shot setting with ResNet-50 as the backbone. This enhancement highlights the superior capability of our model to leverage the integrated multimodal features effectively. Additionally, MIRROR consistently outperforms TANGLE across all settings, further emphasizing its robust feature representation and generalizability. These results highlight the effectiveness of MIRROR in tackling complex survival prediction tasks, reinforcing its potential for improving prognostic modeling in computational pathology.
@@ IV-D Ablation Study
TABLE IV: Ablation Study of Model Components on TCGA-BRCA.

### v3

@@ IV-A Datasets
Four distinct cohorts from the publicly available TCGA dataset [17] were utilized for evaluation: Breast Invasive Carcinoma (BRCA), Non-Small Cell Lung Cancer (NSCLC), Renal Cell Carcinoma (RCC), and Colon and Rectal Adenocarcinoma (COADREAD). Each cohort comprises specific subtypes, as detailed in Table II.
@@
WSIs obtained from these cohorts were processed into patches at $20\times$ magnification, and feature representations were extracted using ResNet-50 [60] and Phikon [61]. Corresponding transcriptomics data were collected from Xena [62] and preprocessed using the proposed novel pipeline, reducing the original 198,620 genes to 40,146 for the BRCA cohort, 10,234 for the NSCLC cohort, 10,303 for the RCC cohort, and 20,056 for the COADREAD cohort, forming the novel refined transcriptomics datasets.
TABLE III: Cancer subtype classification on TCGA-BRCA, TCGA-NSCLC, TCGA-RCC, and TCGA-COADREAD.
Dataset Backbone Setting ABMIL ILRA-MIL PORPOISE SurvPath Linear Classifier TANGLE MIRROR
Acc F1 Acc F1 Acc F1 Acc F1 Acc F1 Acc F1 Acc F1
TCGA-
NSCLC
ResNet-50 10-shot 0.633 0.572 0.634 0.596 0.742 0.699 0.654 0.618 0.660 0.658 0.909 0.909 0.930 0.930
($\pm 0.084$) ($\pm 0.147$) ($\pm 0.072$) ($\pm 0.146$) ($\pm 0.182$) ($\pm 0.242$) ($\pm 0.049$) ($\pm 0.114$) ($\pm 0.024$) ($\pm 0.025$) ($\pm 0.033$) ($\pm 0.033$) ($\pm 0.006$) ($\pm 0.006$)
All Data 0.870 0.868 0.890 0.886 0.986 0.986 0.967 0.966 0.845 0.843 0.989 0.987 0.992 0.992
($\pm 0.015$) ($\pm 0.016$) ($\pm 0.014$) ($\pm 0.015$) ($\pm 0.007$) ($\pm 0.007$) ($\pm 0.027$) ($\pm 0.027$) ($\pm 0.014$) ($\pm 0.015$) ($\pm 0.009$) ($\pm 0.009$) ($\pm 0.011$) ($\pm 0.011$)
Phikon 10-shot 0.709 0.703 0.682 0.678 0.786 0.767 0.739 0.670 0.699 0.694 0.925 0.924 0.939 0.938
($\pm 0.030$) ($\pm 0.029$) ($\pm 0.043$) ($\pm 0.046$) ($\pm 0.155$) ($\pm 0.183$) ($\pm 0.060$) ($\pm 0.079$) ($\pm 0.030$) ($\pm 0.029$) ($\pm 0.036$) ($\pm 0.039$) ($\pm 0.017$) ($\pm 0.017$)
All Data 0.906 0.905 0.917 0.916 0.984 0.992 0.907 0.946 0.901 0.900 0.982 0.982 0.994 0.994
($\pm 0.010$) ($\pm 0.010$) ($\pm 0.017$) ($\pm 0.017$) ($\pm 0.007$) ($\pm 0.008$) ($\pm 0.023$) ($\pm 0.022$) ($\pm 0.008$) ($\pm 0.008$) ($\pm 0.012$) ($\pm 0.012$) ($\pm 0.007$) ($\pm 0.007$)
TCGA-
BRCA
ResNet-50 10-shot 0.796 0.480 0.714 0.402 0.902 0.726 0.860 0.474 0.862 0.507 0.910 0.757 0.910 0.786
($\pm 0.075$) ($\pm 0.028$) ($\pm 0.322$) ($\pm 0.157$) ($\pm 0.010$) ($\pm 0.152$) ($\pm 0.031$) ($\pm 0.014$) ($\pm 0.031$) ($\pm 0.076$) ($\pm 0.036$) ($\pm 0.121$) ($\pm 0.032$) ($\pm 0.092$)
All Data 0.860 0.462 0.926 0.831 0.923 0.894 0.911 0.862 0.901 0.762 0.956 0.909 0.958 0.902
($\pm 0.030$) ($\pm 0.001$) ($\pm 0.016$) ($\pm 0.035$) ($\pm 0.011$) ($\pm 0.022$) ($\pm 0.030$) ($\pm 0.008$) ($\pm 0.029$) ($\pm 0.020$) ($\pm 0.018$) ($\pm 0.029$) ($\pm 0.017$) ($\pm 0.039$)
Phikon 10-shot 0.712 0.589 0.807 0.566 0.902 0.731 0.876 0.671 0.758 0.613 0.921 0.829 0.923 0.832
($\pm 0.131$) ($\pm 0.078$) ($\pm 0.040$) ($\pm 0.060$) ($\pm 0.030$) ($\pm 0.155$) ($\pm 0.111$) ($\pm 0.074$) ($\pm 0.059$) ($\pm 0.065$) ($\pm 0.031$) ($\pm 0.068$) ($\pm 0.025$) ($\pm 0.051$)
All Data 0.919 0.825 0.926 0.848 0.934 0.899 0.890 0.862 0.924 0.821 0.952 0.887 0.955 0.902
($\pm 0.008$) ($\pm 0.021$) ($\pm 0.025$) ($\pm 0.030$) ($\pm 0.030$) ($\pm 0.009$) ($\pm 0.030$) ($\pm 0.009$) ($\pm 0.038$) ($\pm 0.092$) ($\pm 0.022$) ($\pm 0.063$) ($\pm 0.023$) ($\pm 0.047$)
TCGA-
RCC
ResNet-50 10-shot 0.752 0.629 0.769 0.659 0.937 0.846 0.801 0.813 0.752 0.673 0.938 0.832 0.942 0.853
($\pm 0.059$) ($\pm 0.121$) ($\pm 0.020$) ($\pm 0.104$) ($\pm 0.043$) ($\pm 0.112$) ($\pm 0.042$) ($\pm 0.090$) ($\pm 0.016$) ($\pm 0.099$) ($\pm 0.034$) ($\pm 0.111$) ($\pm 0.028$) ($\pm 0.126$)
All Data 0.908 0.677 0.954 0.872 0.976 0.857 0.943 0.907 0.933 0.827 0.988 0.903 0.998 0.932
($\pm 0.017$) ($\pm 0.144$) ($\pm 0.011$) ($\pm 0.139$) ($\pm 0.031$) ($\pm 0.194$) ($\pm 0.062$) ($\pm 0.186$) ($\pm 0.007$) ($\pm 0.116$) ($\pm 0.007$) ($\pm 0.138$) ($\pm 0.003$) ($\pm 0.150$)
Phikon 10-shot 0.839 0.733 0.845 0.751 0.951 0.854 0.924 0.832 0.828 0.727 0.950 0.867 0.952 0.931
($\pm 0.029$) ($\pm 0.083$) ($\pm 0.021$) ($\pm 0.090$) ($\pm 0.030$) ($\pm 0.114$) ($\pm 0.110$) ($\pm 0.061$) ($\pm 0.027$) ($\pm 0.107$) ($\pm 0.014$) ($\pm 0.123$) ($\pm 0.030$) ($\pm 0.042$)
All Data 0.968 0.886 0.976 0.902 0.988 0.920 0.977 0.873 0.961 0.885 0.989 0.923 0.997 0.931
($\pm 0.015$) ($\pm 0.133$) ($\pm 0.007$) ($\pm 0.016$) ($\pm 0.011$) ($\pm 0.015$) ($\pm 0.029$) ($\pm 0.027$) ($\pm 0.011$) ($\pm 0.130$) ($\pm 0.001$) ($\pm 0.147$) ($\pm 0.003$) ($\pm 0.149$)
TCGA-
COADREAD
ResNet-50 10-shot 0.815 0.542 0.781 0.461 0.813 0.448 0.820 0.476 0.819 0.502 0.831 0.539 0.834 0.556
($\pm 0.057$) ($\pm 0.066$) ($\pm 0.102$) ($\pm 0.034$) ($\pm 0.065$) ($\pm 0.020$) ($\pm 0.055$) ($\pm 0.049$) ($\pm 0.055$) ($\pm 0.080$) ($\pm 0.043$) ($\pm 0.066$) ($\pm 0.039$) ($\pm 0.144$)
All Data 0.828 0.515 0.813 0.448 0.814 0.447 0.830 0.523 0.816 0.470 0.928 0.881 0.938 0.893
($\pm 0.051$) ($\pm 0.091$) ($\pm 0.065$) ($\pm 0.020$) ($\pm 0.065$) ($\pm 0.020$) ($\pm 0.065$) ($\pm 0.020$) ($\pm 0.069$) ($\pm 0.064$) ($\pm 0.042$) ($\pm 0.051$) ($\pm 0.022$) ($\pm 0.017$)
Phikon 10-shot 0.659 0.495 0.691 0.512 0.813 0.448 0.812 0.446 0.644 0.510 0.772 0.483 0.828 0.515
($\pm 0.133$) ($\pm 0.099$) ($\pm 0.102$) ($\pm 0.068$) ($\pm 0.065$) ($\pm 0.020$) ($\pm 0.065$) ($\pm 0.020$) ($\pm 0.118$) ($\pm 0.073$) ($\pm 0.087$) ($\pm 0.103$) ($\pm 0.043$) ($\pm 0.101$)
All Data 0.834 0.554 0.813 0.548 0.859 0.552 0.815 0.449 0.819 0.570 0.928 0.880 0.941 0.902
($\pm 0.037$) ($\pm 0.114$) ($\pm 0.065$) ($\pm 0.020$) ($\pm 0.065$) ($\pm 0.217$) ($\pm 0.065$) ($\pm 0.020$) ($\pm 0.069$) ($\pm 0.096$) ($\pm 0.049$) ($\pm 0.046$) ($\pm 0.020$) ($\pm 0.018$)
@@ IV-B1 Data Preprocessing
For WSIs, adhering to the conventional preprocessing protocol established in [63], the foreground tissue regions are first segmented using Otsu’s method [64]. These segmented regions are then divided into patches, forming an instance bag that is processed by a pre-trained patch encoder to extract feature representations. To enable batched training, we perform random sampling to select a fixed number of feature representations, employing replacement if the available number of representations is less than the required count.
@@ IV-B2 Experiment Setup
All experiments, including both pretraining and downstream task evaluations, are conducted under a 5-fold cross-validation protocol. Each dataset is randomly partitioned into five folds. In each run, one fold is held out for evaluation while the remaining four folds are used for training. Specifically, under the 10-shot setting, training samples are randomly selected from the four training folds. The evaluation fold remains strictly held out and is never accessed during training, ensuring no data leakage. All split definitions and sampling scripts are released in our public code repository to support reproducibility.
Pretraining is performed separately for each cohort and repeated under each cross-validation split. This ensures that no evaluation samples are used during representation learning, maintaining a fair and unbiased evaluation setting.
After pretraining, MIRROR is further evaluated on cancer subtyping and survival analysis tasks by concatenating the outputs from the two modality-specific encoders and feeding the combined vector into a linear classifier. We adopt both few-shot learning and linear probing for evaluation, following the same 5-fold cross-validation setup. To further assess generalization ability, we conduct evaluations using two different pretrained patch encoders.
The model is optimized using the Adam optimizer [65] and is implemented using PyTorch [66]. All training sessions are performed on a single NVIDIA GeForce RTX 3090 GPU paired with an Intel i7-12700K CPU and 32GB of memory. Full hyperparameter configurations for both pretraining and downstream task evaluations are provided in Table I. All hyperparameter settings and training configurations are also released in our public code repository.
@@ IV-C Downstream Tasks
Two tasks are used for downstream evaluations, namely subtype classification and survival analysis. We compared the proposed MIRROR against ABMIL [67], ILRA-MIL [68], PORPOISE [69], SurvPath [14], Linear Classifier, and TANGLE [12], utilizing two distinct backbones under both 10-shot and linear probing settings. Among these methods, ABMIL, ILRA-MIL, and Linear Classifier take histopathology as input, whereas the other models integrate both histopathology and transcriptomics data. To ensure fair comparisons and optimal performance, all methods are trained on the proposed transcriptomic datasets. Notably, the original TANGLE framework evaluates only histopathology features. To comprehensively assess its performance and maintain methodological consistency, we extend TANGLE by incorporating both histopathology and transcriptomic features, aligning its input with that of MIRROR.
@@ IV-C1 Subtype Classification
As illustrated in Table III, the proposed MIRROR consistently achieves superior performance, demonstrating its effectiveness in integrating histopathology and transcriptomic data. On the TCGA-NSCLC dataset, MIRROR achieves the highest accuracy and F1 score across all settings, outperforming all baseline methods. Notably, its strong performance remains consistent regardless of the chosen backbone architecture, highlighting its robustness on the TCGA-NSCLC dataset. Similarly, on the TCGA-BRCA dataset, MIRROR achieves state-of-the-art results, surpassing the second best model’s (TANGLE) result by 1.5% in F1 score under the linear probing setting with the Phikon backbone. Additionally, MIRROR significantly outperforms TANGLE by 6.4% in F1 score on the TCGA-RCC dataset under the 10-shot setting with the Phikon backbone. On the TCGA-COADREAD dataset, it surpasses TANGLE by 1.3% in accuracy under the linear probing setting with the Phikon backbone. These findings validate the robustness and generalizability of MIRROR across different cancer subtypes and backbone architectures.
TABLE IV: Cancer survival analysis on TCGA-BRCA and TCGA-NSCLC, TCGA-RCC, and TCGA-COADREAD.
Dataset Backbone Setting ABMIL ILRA-MIL PORPOISE SurvPath Linear Classifier TANGLE MIRROR
TCGA-
NSCLC
ResNet-50 10-shot 0.529 ($\pm 0.033$) 0.593 ($\pm 0.045$) 0.565 ($\pm 0.011$) 0.547 ($\pm 0.071$) 0.604 ($\pm 0.019$) 0.570 ($\pm 0.055$) 0.605 ($\pm 0.011$)
All Data 0.538 ($\pm 0.042$) 0.606 ($\pm 0.038$) 0.614 ($\pm 0.047$) 0.613 ($\pm 0.048$) 0.618 ($\pm 0.041$) 0.565 ($\pm 0.042$) 0.621 ($\pm 0.054$)
Phikon 10-shot 0.557 ($\pm 0.016$) 0.574 ($\pm 0.008$) 0.577 ($\pm 0.022$) 0.573 ($\pm 0.056$) 0.577 ($\pm 0.017$) 0.583 ($\pm 0.049$) 0.584 ($\pm 0.007$)
All Data 0.567 ($\pm 0.039$) 0.608 ($\pm 0.025$) 0.602 ($\pm 0.053$) 0.572 ($\pm 0.031$) 0.596 ($\pm 0.035$) 0.593 ($\pm 0.039$) 0.613 ($\pm 0.043$)
TCGA-
BRCA
ResNet-50 10-shot 0.575 ($\pm 0.054$) 0.587 ($\pm 0.041$) 0.601 ($\pm 0.028$) 0.560 ($\pm 0.039$) 0.605 ($\pm 0.044$) 0.607 ($\pm 0.052$) 0.612 ($\pm 0.046$)
All Data 0.573 ($\pm 0.077$) 0.616 ($\pm 0.026$) 0.659 ($\pm 0.091$) 0.651 ($\pm 0.019$) 0.657 ($\pm 0.029$) 0.580 ($\pm 0.033$) 0.671 ($\pm 0.096$)
Phikon 10-shot 0.544 ($\pm 0.027$) 0.616 ($\pm 0.067$) 0.608 ($\pm 0.031$) 0.603 ($\pm 0.044$) 0.587 ($\pm 0.042$) 0.602 ($\pm 0.039$) 0.623 ($\pm 0.054$)
All Data 0.551 ($\pm 0.029$) 0.645 ($\pm 0.065$) 0.500 ($\pm 0.068$) 0.650 ($\pm 0.072$) 0.658 ($\pm 0.048$) 0.540 ($\pm 0.055$) 0.665 ($\pm 0.082$)
TCGA-
RCC
ResNet-50 10-shot 0.598 ($\pm 0.047$) 0.647 ($\pm 0.044$) 0.669 ($\pm 0.046$) 0.641 ($\pm 0.040$) 0.651 ($\pm 0.051$) 0.603 ($\pm 0.072$) 0.697 ($\pm 0.013$)
All Data 0.596 ($\pm 0.030$) 0.754 ($\pm 0.048$) 0.777 ($\pm 0.020$) 0.750 ($\pm 0.040$) 0.743 ($\pm 0.040$) 0.794 ($\pm 0.040$) 0.800 ($\pm 0.045$)
Phikon 10-shot 0.503 ($\pm 0.040$) 0.672 ($\pm 0.041$) 0.694 ($\pm 0.022$) 0.630 ($\pm 0.058$) 0.651 ($\pm 0.056$) 0.677 ($\pm 0.083$) 0.739 ($\pm 0.032$)
All Data 0.502 ($\pm 0.041$) 0.783 ($\pm 0.046$) 0.780 ($\pm 0.039$) 0.756 ($\pm 0.037$) 0.778 ($\pm 0.051$) 0.801 ($\pm 0.044$) 0.803 ($\pm 0.043$)
TCGA-
COADREAD
ResNet-50 10-shot 0.591 ($\pm 0.048$) 0.569 ($\pm 0.030$) 0.598 ($\pm 0.042$) 0.571 ($\pm 0.063$) 0.595 ($\pm 0.065$) 0.573 ($\pm 0.071$) 0.619 ($\pm 0.050$)
All Data 0.601 ($\pm 0.050$) 0.631 ($\pm 0.057$) 0.616 ($\pm 0.054$) 0.671 ($\pm 0.093$) 0.706 ($\pm 0.054$) 0.574 ($\pm 0.071$) 0.730 ($\pm 0.057$)
Phikon 10-shot 0.490 ($\pm 0.054$) 0.611 ($\pm 0.075$) 0.601 ($\pm 0.112$) 0.620 ($\pm 0.066$) 0.639 ($\pm 0.053$) 0.630 ($\pm 0.094$) 0.657 ($\pm 0.054$)
All Data 0.538 ($\pm 0.021$) 0.713 ($\pm 0.024$) 0.642 ($\pm 0.047$) 0.660 ($\pm 0.068$) 0.701 ($\pm 0.043$) 0.618 ($\pm 0.076$) 0.721 ($\pm 0.033$)
@@ IV-C2 Survival Prediction
As shown in Table IV, the proposed MIRROR achieves significant performance improvements over all baseline methods. On the TCGA-NSCLC dataset, MIRROR surpasses PORPOISE by $4.0\%$ and achieves a $7.6\%$ improvement compared to ABMIL under the 10-shot setting with ResNet-50 as the backbone. This enhancement highlights the superior capability of our model to leverage the integrated multimodal features effectively. Additionally, MIRROR consistently outperforms TANGLE across all settings, further emphasizing its robust feature representation and generalizability. These results highlight the effectiveness of MIRROR in tackling complex survival prediction tasks, reinforcing its potential for improving prognostic modeling in computational pathology.
@@ IV-D Ablation Study
IV-D1 Ablation Study of Model Components
We conducted an ablation study to evaluate the effectiveness of the key components of MIRROR, including the modality alignment module, the modality retention module, and the style clustering module. The study was performed on the TCGA-NSCLC, TCGA-BRCA, TCGA-RCC, and TCGA-COADREAD datasets using linear probing with Phikon as the backbone, comparing MIRROR’s performance both with and without these components.
As shown in Table V, the results highlight the contributions of each module. The modality retention module enhances subtype classification, demonstrating its ability to preserve modality-specific information. However, its inclusion slightly decreases the performance of survival prediction. This reduction arises because the module retains not only disease-relevant features but also disease-irrelevant variations, such as staining artifacts or non-prognostic gene fluctuations. Survival prediction is particularly sensitive to such variance, as censoring and the resulting class imbalance reduce label informativeness and lower the effective signal-to-noise ratio compared with subtype classification. These nuisance variations can distort risk calibration and disrupt the relative ordering of patients, thereby depressing the C-index. The style clustering module effectively addresses this limitation by encouraging the model to emphasize disease-relevant patterns while suppressing redundancy, which yields consistent gains in survival prediction. While its inclusion results in a minor decline in subtype classification on TCGA-BRCA, performance remains stable on other datasets, and this small drawback is fully mitigated when both modules are jointly applied. Ultimately, MIRROR achieves the best overall performance, confirming the synergistic benefits of the proposed components.
IV-D2 Ablation Study of Style Clustering Module
To empirically validate the claims in Section III-E, we evaluated the individual and joint contributions of the style module and the clustering module. In the style ablation, the stochastic encoder was replaced with a deterministic MLP that produced a direct embedding $z$ instead of learning distributional parameters $\mu$ and $\log\sigma$. In the clustering ablation, the prototype-based soft alignment was replaced with a hard alignment between histopathology and transcriptomics embeddings, thereby enforcing a rigid matching strategy in place of probabilistic prototype consistency. All experiments employed Phikon as the backbone under the linear probing setting on TCGA-NSCLC, TCGA-BRCA, TCGA-RCC, and TCGA-COADREAD.
The results presented in Table VI demonstrate that the two modules function as complementary forces that jointly achieve compression and promotion of disease-relevant representations. The style loss enforces compactness by penalizing deviations from the standard normal prior, which suppresses redundant degrees of freedom and reduces variance. When used in isolation, this compression may suppress both disease-relevant and irrelevant information, thereby diminishing overall performance. In contrast, the clustering loss promotes disease-relevant alignment by encouraging convergence of the probabilistic assignments toward shared prototypes. Applied alone, this alignment can also admit nuisance variability, since there is no mechanism to constrain the latent space. When both modules are combined, compression removes redundant variation while alignment preserves and organizes task-relevant factors, resulting in consistent improvements across cohorts and tasks. These findings confirm the intended design of the style clustering module, in which the two objectives counterbalance each other and guide the representations toward a stable state that eliminates redundancy while retaining disease-relevant structure.
TABLE V: Ablation study of model components.
Module Setting TCGA-NSCLC TCGA-BRCA TCGA-RCC TCGA-COADREAD
Alignment Retention Style Clustering Subtyping Survival Subtyping Survival Subtyping Survival Subtyping Survival
0.986 0.600 0.953 0.612 0.992 0.775 0.898 0.697
(± $\pm\,0.009$) (± $\pm\,0.040$) (± $\pm\,0.024$) (± $\pm\,0.075$) (± $\pm\,0.003$) (± $\pm\,0.051$) (± $\pm\,0.023$) (± $\pm\,0.089$)
0.988 0.595 0.954 0.610 0.994 0.773 0.918 0.647
(± $\pm\,0.012$) (± $\pm\,0.024$) (± $\pm\,0.028$) (± $\pm\,0.085$) (± $\pm\,0.003$) (± $\pm\,0.040$) (± $\pm\,0.059$) (± $\pm\,0.049$)
0.989 0.610 0.950 0.654 0.996 0.778 0.921 0.702
(± $\pm\,0.003$) (± $\pm\,0.044$) (± $\pm\,0.016$) (± $\pm\,0.083$) (± $\pm\,0.003$) (± $\pm\,0.045$) (± $\pm\,0.027$) (± $\pm\,0.034$)
0.994 0.613 0.955 0.665 0.997 0.803 0.941 0.721
(± $\pm\,0.007$) (± $\pm\,0.043$) (± $\pm\,0.023$) (± $\pm\,0.082$) (± $\pm\,0.003$) (± $\pm\,0.043$) (± $\pm\,0.020$) (± $\pm\,0.033$)
IV-D3 Ablation Study of Gene Preprocessing Methods
We compared our preprocessing pipeline with two widely used alternatives. The first is the preprocessed transcriptomics data from PORPOISE [69], which collects raw data from cBioPortal [70] and reduces the dimensionality of the collected RNA sequencing features by selecting gene sets from the gene family categories in the Molecular Signatures Database (MSigDB) [71]. The second is the preprocessing adopted by SurvPath [14], where raw transcriptomic profiles are obtained from the Xena [62] and pathways are constructed from Reactome [72] and the MSigDB Hallmarks collection. Although SurvPath operates at the pathway rather than the gene level, we include it here for completeness. Since neither PORPOISE nor SurvPath has released their preprocessing code, and thus our comparison relies on the processed data they provide, with TCGA-BRCA being the only cohort aligned with our setting. All experiments employed Phikon as the backbone under the linear probing setting on TCGA-BRCA. As summarized in Table VII, our preprocessing approach consistently achieves the best overall performance across both subtyping and survival tasks, outperforming PORPOISE and SurvPath. These results confirm the effectiveness of our preprocessing strategy.
IV-D4 Ablation Study of Loss Term Weights
To determine the hyperparameters in Eq. (22), we conducted an ablation study on the loss term weights. The evaluation employed Phikon as the backbone for feature extraction under the linear probing setting on TCGA-BRCA and followed the same 5-fold cross-validation protocol. As reported in Table VIII, we varied the relative contributions of $\lambda_{\alpha}$, $\lambda_{\beta}$, and $\lambda_{\gamma}$, which respectively balance the alignment loss $L_{\text{align}}$, the retention loss $L_{\text{retention}}$, and the style clustering loss $L_{\text{style clustering}}$, while keeping their sum equal to one. The results indicate that performance remains stable across different configurations, with all three settings producing comparable outcomes. The configuration $\lambda_{\alpha}=0.5$, $\lambda_{\beta}=0.3$, and $\lambda_{\gamma}=0.2$ provided the best overall balance, delivering the highest accuracy for subtype classification and the strongest C-index for survival prediction. Based on these findings, we adopt this configuration throughout the manuscript. This analysis demonstrates that our method is not overly sensitive to the precise choice of hyperparameters and that the reported improvements reflect robust performance rather than fragile tuning.
IV-E Qualitative Analysis
Here, we demonstrate the explainability of MIRROR with attention weights visualization from the slide encoder and UMAP [73] projections of features extracted by slide and RNA encoders.
As illustrated in Figure 4, we selected four representative samples from each cohort used. The attention weights from the slide encoder are visualized using heatmaps to highlight regions of interest. Areas shown in warmer colors indicate regions that the model pays more attention to. The results suggest MIRROR consistently attends to disease-relevant regions, particularly to cancerous regions. This strong alignment showcases MIRROR’s extraordinary interpretability, confirming that it not only excels in downstream tasks but also provides biologically grounded decision-making insights. Across all four TCGA cohorts, regions receiving higher attention consistently correspond to cancerous tissue characterized by dense malignant cellularity and invasive fronts, whereas regions with lower attention are predominantly associated with non-neoplastic compartments, including background stroma, adipose tissue, and non-tissue artifacts. These consistent patterns offer a clear biological interpretation of both high- and low-attention patches.
TABLE VI: Ablation study of style clustering module.
Module Setting TCGA-NSCLC TCGA-BRCA TCGA-RCC TCGA-COADREAD
Style Clustering Subtyping Survival Subtyping Survival Subtyping Survival Subtyping Survival
0.984 0.596 0.939 0.625 0.993 0.772 0.887 0.636
(± $\pm\,0.007$) (± $\pm\,0.023$) (± $\pm\,0.026$) (± $\pm\,0.035$) (± $\pm\,0.004$) (± $\pm\,0.051$) (± $\pm\,0.060$) (± $\pm\,0.039$)
0.985 0.598 0.945 0.634 0.994 0.768 0.896 0.633
(± $\pm\,0.006$) (± $\pm\,0.033$) (± $\pm\,0.021$) (± $\pm\,0.044$) (± $\pm\,0.004$) (± $\pm\,0.033$) (± $\pm\,0.024$) (± $\pm\,0.048$)
0.994 0.613 0.955 0.665 0.997 0.803 0.941 0.721
(± $\pm\,0.007$) (± $\pm\,0.043$) (± $\pm\,0.023$) (± $\pm\,0.082$) (± $\pm\,0.003$) (± $\pm\,0.043$) (± $\pm\,0.020$) (± $\pm\,0.033$)
TABLE VII: Ablation study of gene preprocessing methods.
Preprocessing TANGLE MIRROR
Subtyping Survival Subtyping Survival
PORPOISE 0.929 0.534 0.931 0.656
(± $\pm\,0.034$) (± $\pm\,0.099$) (± $\pm\,0.028$) (± $\pm\,0.076$)
SurvPath 0.931 0.535 0.938 0.628
(± $\pm\,0.014$) (± $\pm\,0.088$) (± $\pm\,0.008$) (± $\pm\,0.079$)
MIRROR 0.952 0.540 0.955 0.665
(± $\pm\,0.022$) (± $\pm\,0.055$) (± $\pm\,0.023$) (± $\pm\,0.082$)
TABLE VIII: Ablation study of loss term weights.